# Stochastic trade-offs and the emergence of diversification in E. coli evolution experiments

### v2

@@ III.1 Reproducibility and variability of experimental results
Figure 2: Reproducibility and variability of evolutionary branching in our mesoscopic model with a stochastic trade-off. a-c) Sub-populations in phenotypic space (represented by a red circle whose radius is proportional to its abundance) for three different times along the evolutionary trajectory, pre-branching (a), during branching (b) and after branching (c), for one specific realization with $\sigma=4\cdot 10^{-4}$. The inset in b) is a zoom in the region close to the branching point (yellow star), where $d\approx 10^{-5}$ is the maximum distance from a subpopulation to the average of the stochastic trade-off. d) Percentage of realizations (number of them $N_{r}=200$) that undergo evolutionary branching for different amplitudes of the stochastic trade-off. e-g) Probability density distributions across experiments (number of realizations $N_{r}=500$) of: branching times for different stochastic trade-off amplitudes (e); and fraction of acetate scavengers at different times for $\sigma=0$ (f) and $\sigma=3\cdot 10^{-4}$ (g) (see Methods Sec. VI.4). System’s parameters for all the plots are: $\beta=0.25$ and those specified in Table 1 where the shown $V_{g/a}$ are rescaled with respect to the table value as $V_{a/g}=\gamma V_{a/g}^{\text{table}}$ with $\gamma=4.5\cdot 10^{10}$. Stochastic trade-off parameters are: $(a_{tr},b_{tr},c_{tr},\tau)=(83.3,-3.5,0.03,10)$ (see Methods Sec.VI.5 for definitions and further information). Evolutionary simulations are carried out as described in Methods Sec.VI.3.
@@
As illustrated in Figure 2, computational analyses of the eco-evolutionary (mesoscopic) model coupled with a stochastic trade-off (as defined in Methods Sec.VI.2) reveal that it is able to account for evolutionary branching as well as for the variability evinced in the experiments [16]. In particular, panels (a-c) show three temporal snapshots of the evolution of the population in phenotypic space for a specific realization of the model with the stochastic trade-off (as illustrated in the inset of panel b) that ended up diversifying in two highly-specialized sub-populations.
@@
Panel (a) represents the initial phase of the evolutionary picture, in which a glucose-specialist population that produces acetate evolves its ability to assimilate acetate because it is evolutionarily-favoured so that it moves progressively “uphill”. Panel (b) shows the phase in which the whole population –upon arriving to a singular point— splits in two main ecotypes due to disruptive selection. Panel (c) illustrates the last phase in which each of the two branched ecotypes specializes more and more in metabolizing preferentially one of the two carbon sources, thus giving rise to the two different ecotypes reported in the experiments —the glucose specialist and the acetate scavenger. This evolutionary picture is compatible with the one described in the experiments [15].
@@
On the other hand, panel (d) summarizes the amount of realizations that, starting from a glucose-specialist population and with the same parameters as in panel (a), ended up undergoing evolutionary branching at a maximum evolutionary time (arbitrarily fixed to $t_{\text{evo}}^{\text{max}}=3\cdot 10^{4}$) for different amplitudes of the stochastic trade-off region, as obtained by varying the noise parameter, $\sigma$, for constant $\tau$. In this way, by tuning the noise amplitude —or equivalently the flexibility of the metabolic constraints— the model is able to account for the experimental fact that diversification is observed in variable proportion of the experiments.
@@
Furthermore, as shown in panels (e-g) the model can also explain the empirically-observed variability in both the branching times and the proportion of acetate scavengers after population splitting. Indeed, panel (e) reveals that branching times —estimated as explained in Methods Sec.VI.4— are variable and their distribution depends on the noise amplitude, $\sigma$. One can observe that, in the case with a ”hard” trade-off (i.e. $\sigma=0$), branching occurs almost always at the same time (strongly peaked distribution), but the larger the noise $\sigma$ (or equivalently, the amplitude of the trade-off), the wider the distributions of branching times. On the other hand, Fig.2f shows how the distribution of the fraction of acetate scavengers at different times is quite peaked for the case with the hard trade-off, whereas the case with a non-vanishing amplitude of the trade-off, Fig.2g, shows a much wider distribution.
@@
Let us remark that, as argued before, the finite-widths distributions that emerge in the absence of stochasticity in the trade-off, i..e for $\sigma=0$, stem from the finite time we allow for the ecological dynamics to stabilize, which may result in the presence of clonal interference, and the intrinsic randomness of the evolutionary algorithm (see Methods Sec.VI.3). We have verified that the distributions progressively thin out as the maximum time allowed for the ecological dynamics to relax is extended (data not shown).
@@ III.2 Branching feasibility and trade-off (in)dependence
Figure 3: Branching feasibility and the fraction of realizations that undergo branching. a) Branching feasibility plot: contour plot that illustrates in phenotypic space the range of local curvatures of the hard trade-off that allow evolutionary branching. This curvature range can also be interpreted (with the stochastic trade-off idea) as the likeliness of having evolutionary branching at each point. b-c) Percentage of realizations (total number, $N_{r}=200$) of the eco-evolutionary model with the stochastic trade-off that undergo evolutionary branching for two different maximum evolutionary times $t_{\text{evo}}^{\text{max}}=2.5\cdot 10^{4}$ and $t_{\text{evo}}^{\text{max}}=5\cdot 10^{4}$ at different points in phenotypic space (that have different curvature range or equivalently different branching feasibility). Points in (b) are those in yellow in panel (a) and $\sigma=10^{-4}$. Points in (c) are those in bourdeaux in panel (a) and $\sigma=3.5\cdot 10^{-4}$. System’s parameters are: $\beta=0.25$ and those specified in Table 1 where the shown $V_{g/a}$ are rescaled with respect to the table value as $V_{a/g}=\gamma V_{a/g}^{\text{table}}$ with $\gamma=4.5\cdot 10^{10}$. Stochastic trade-off parameters are: $\tau=10$ and the others are obtained for each point shown in panel a) as specified in Methods Sec.VI.5 with $\delta=0.25$. Eco-evolutionary simulations are carried out as described in Methods Sec.VI.3.
@@
To this aim, we make use of the formalism proposed by de Mazancourt and Dieckmann [44]. As originally developed, such a framework allows one to analyse geometrically the possible evolutionary outcomes that can occur in the deterministic setting without imposing a specific mathematical function for the trade-off (for the sake of completeness, an overview of the method is described in Methods Sec.VI.1). In particular, within this formalism one can construct plots such as that of Fig. 3a, providing information on how independent is evolutionary branching on the specific mathematical form of the hard trade-off. More specifically, it gives at each point in phenotypic space, a measure of the range of possible local curvatures of the hard trade-off function that allows the model to undergo evolutionary branching at such a point, thus quantifying the robustness of branching against changes in the shape of the trade-off function. Regions where the range of possible local curvatures is small (deep blue in Fig. 3a), allow the model to undergo evolutionary branching deterministically only for very specific (fine-tuned) trade-off curves. Instead, in regions where the curvature range is high (light green in Fig. 3a), evolutionary branching occurs deterministically for a larger family of possible trade-off functions (see Section VI.1 for further details).
Now, we would like to see how this (deterministic) formalism could be reinterpreted when we include the idea of stochastic trade-offs. It turns out that using stochastic trade-off, the previous plots (e.g. Fig.3a) can be interpreted as “evolutionary-branching feasibility plots”, i.e. graphs showing the regions in phenotypic space where it is more likely to see an evolutionary branching event. In other words, regions in phenotypic space where branching is more robust against changes in the specific form of the hard trade-off correspond to regions in phenotypic space where branching is more likely to happen with the stochastic trade-off (stochastic) trade-off and vice-versa.
To justify this claim, we employ without loss of generality our eco-evolutionary model with a certain stochastic trade-off, whose average curve produces branching deterministically (i.e. in the $\sigma=0$ limit) at a selected point in phenotypic space (see Methods Sec.VI.2 for further details on how the stochastic trade-off is built). Then, as illustrated in Fig.3b-c, points where deterministic branching depends weakly on the specific form of the hard trade-off —i.e. points where the range of local curvatures of the hard trade-off that allow branching is larger (brighter colours in Fig.3a)— correspond in the stochastic trade-off case to more realizations that end up experiencing diversification at that point and vice-versa.
@@
Therefore, interpreting Fig.3a as a branching-feasibility plot, we observe that branching is more likely for low values of $V_{a}$, which agrees with the reported picture of the evolution experiments given in [15]. As shown in Sec.III.3, the specific contour pattern depends highly on the conditions of the experiment, that is, the concentration of glucose in the environment supplying the chemostat, $G_{0}$, the dilution rate, $d$, and the proportion of the consumed glucose that gets excreted as acetate, $\beta$. Thus, Fig.3a reveals essentially that the system, under these specific environmental conditions, is rather likely to undergo evolutionary branching, which justifies its observed repeatability in actual experiments.
@@ III.3 Theoretical predictions for different experimental values
Figure 4: Branching feasibility and the percentage of realizations that undergo branching for different experimental conditions. a-c) Branching-feasibility plots for different values of glucose concentration in the environment supplying the chemostat: $G_{0}=260$ (a), $G_{0}=694$ (b) and $G_{0}=3469$ (c), all of them in $\mu$mol/L. d) Number of realizations that undergo branching at the point depicted in a-c) for the same stochastic trade-off with $(a_{tr},b_{tr},c_{tr},\tau,\sigma)=(88.2,-2.5,0.014,10,6\cdot 10^{-4})$ (see ’Stochastic trade-offs’ in Materials and Methods for definitions and further information). e-g) Branching-feasibility plots for different values of acetate concentration in the environment supplying the chemostat: $A_{0}=100$ (e), $A_{0}=260$ (f) and $A_{0}=1000$ (g), all of them in $\mu$mol/L. h) Number of realizations that undergo branching at the point depicted in e-g) for the same stochastic trade-off with $(a_{tr},b_{tr},c_{tr},\tau,\sigma)=(164.3,-3.6,0.011,10,5\cdot 10^{-4})$. i-k) Branching-feasibility plots for different values of the dilution rate: $D=0.0015$ (i), $D=0.006$ (j) and $D=0.015$ (k), all of them in 1/min. l) Number of realizations that undergo branching at the point depicted in i-k) for the same stochastic trade-off with $(a_{tr},b_{tr},c_{tr},\tau,\sigma)=(65.5,-5.6,0.088,10,1.2\cdot 10^{-4})$. Total number of realizations for the histograms is $N_{r}=200$. The remaining parameters specified in Table S1 in the supplementary material where the shown $V_{g/a}$ are rescaled with respect to the table value as $V_{a/g}=\gamma V_{a/g}^{\text{table}}$ with $\gamma=4.5\cdot 10^{10}$.
@@
In particular, Fig.4a-c clearly illustrates that higher concentrations of glucose in the environment supplying the chemostat ($G_{0}$) generally increase the likelihood of branching. This is further verified in Fig.4d. Similar to Fig.3b-c, we use a certain stochastic trade-off for the three different environments whose average curve induces deterministic branching at a selected point. As evident in the histogram, the higher the glucose supply, the greater the number of realizations that undergo evolutionary branching. Moreover, the similarity in the contour pattern shape to that in Fig.3a is attributed to the fact that the proportion of available acetate (i.e. $\beta$ in this case) remains the same.
@@
Finally, it becomes clear from Fig.4i-k that the dilution rate ($D$) is also a crucial parameter for the emergence of branching. Specifically, higher values of $D$ significantly reduce the likelihood of branching, as illustrated in Fig. 4l.
@@ IV Discussion
Previous work [31] showed, in the context of adaptive dynamics, that a simple eco-evolutionary model can explain in a parsimonious way such a diversification event as evolutionary branching, by assuming a specific (hard) trade-off function between the evolving phenotypes (i.e. glucose and acetate). Nevertheless, such an approach is deterministic in nature, i.e. either predicts that branching should be observed always or never depending on the specific trade-off form chosen. Therefore, it is not appropiate to explain the empirically observed stochasticity in the experiments, that is, the frequent emergence of diversification and the variability both in branching times and in the relative weights of the emerging populations.
Alternatively, in this work, we propose the idea of a stochastic trade-off that stands as an effective representation of the underlying metabolic relations that constrain the evolution of bacteria in a high-dimensional phenotypic space. Thus, instead of forcing the evolutionary trajectories in a lower-dimensional phenotypic space to a certain fixed curve, it allows the populations to move in a wider region in such a space, introducing a new source of stochasticity along such a space. The stochasticity of mutations in a high-dimensional feature space translates into variability that is not strictly constrained to a one-dimensional curve in the space of preferences, but rather to a flexible tradeoff curve. Mathematically, the stochastic trade-off is characterized by a central curve that stands as the average of the different trajectories, the correlation time, $\tau$, that defines how fast the trajectories go back to this average curve and its amplitude, which, for fixed $\tau$, is determined by $\sigma$ and defines the width of the (fuzzy) region where populations can evolve.
Thus, we have built a biologically plausible eco-evolutionary model (see Eqs.1-3), fed with experimentally measured parameters (see Table 1 in Materials and Methods) and equipped with the mentioned stochastic trade-off. We have shown that such a model is indeed able to account for the variability reported in the experiments (Fig.2). In particular, contrary to existing models it does reproduce: (i) the fact that branching occurs in a large proportion, but not all, of the realizations (Fig.2a-d) (ii) it may happen at considerably different times (Fig.2f) and (iii) the fraction of acetate scavengers at a certain time are variable across experiments (Fig.2g-h). Let us remark that, some recent studies [45, 46] have also demonstrated variability in the way in which branching occurs across replicated experiments, but the considered ecological models differ from those used in experiments, and, unlike our model, the degree of variability is not easily controllable through a specific parameter.

### v3

@@ III.1 Reproducibility and variability of experimental results
Figure 2: Reproducibility and variability of evolutionary branching in our mesoscopic model with a stochastic trade-off. (a-c) Sub-populations in phenotypic space (represented by a red circle whose radius is proportional to its abundance) for three different times along the evolutionary trajectory, pre-branching (a), during branching (b) and after branching (c), for one specific realization with $\sigma=4\cdot 10^{-4}$. The inset in b) is a zoom in the region close to the branching point (yellow star), where $d\approx 10^{-5}$ is the maximum distance from a subpopulation to the average of the stochastic trade-off. (d) Percentage of realizations (number of them $N_{r}=200$) that undergo evolutionary branching for different amplitudes of the stochastic trade-off. (e-g) Probability density distributions across experiments (number of realizations $N_{r}=500$) of: branching times for different stochastic trade-off amplitudes (with dashed lines highlighting the shift in the peaks of the distributions) (e); and fraction of acetate scavengers at different times for $\sigma=0$ (f) and $\sigma=3\cdot 10^{-4}$ (g) (see SM Sec. IV). System’s parameters for all the plots are: $\beta=0.25$ and those specified in SM Table I where the shown $V_{g/a}$ are rescaled with respect to the table value as $V_{a/g}=\gamma V_{a/g}^{\text{table}}$ with $\gamma=4.5\cdot 10^{10}$. Stochastic trade-off parameters are: $(a_{tr},b_{tr},c_{tr},\tau)=(83.3,-3.5,0.03,10)$ (see SM Sec. V for definitions and further information). Evolutionary simulations are carried out as described in SM Sec. III.
@@
As illustrated in Fig.2, computational analyses of the eco-evolutionary (mesoscopic) model coupled with a stochastic trade-off (as defined in SM Sec. II) reveal that it is able to account for evolutionary branching as well as for the variability evinced in the experiments [16]. In particular, Fig.2(a-c) shows three temporal snapshots of the evolution of the population in phenotypic space for a specific realization of the model with the stochastic trade-off (as illustrated in the inset of panel b) that ended up diversifying in two highly-specialized sub-populations.
@@
Panel (a) represents the initial phase of the evolutionary picture, in which a glucose-specialist population that produces acetate evolves its ability to assimilate acetate because it is evolutionarily-favoured, so that it starts moving progressively “uphill”. Panel (b) shows the phase in which the population has accumulated some variability and –upon arriving to a singular point— splits in two main ecotypes due to disruptive selection: one more specialized in acetate and the other still relying predominantly on glucose consumption. Panel (c) illustrates the final phase, where the two branched types diverge, each becoming increasingly specialized in metabolizing one of the two carbon sources. This divergence thus gives rise to the two distinct ecotypes observed in the experiments: the glucose specialist and the acetate scavenger.
This evolutionary picture is fully compatible with the one described in the experiments [15].
@@
On the other hand, Fig.2d summarizes the amount of realizations that, starting from a glucose-specialist population and with the same parameters as in panel (a), ended up undergoing evolutionary branching at a maximum evolutionary time (arbitrarily fixed to $t_{\text{evo}}^{\text{max}}=3\cdot 10^{4}$) for different amplitudes of the stochastic trade-off region, as obtained by varying the noise parameter, $\sigma$, for constant $\tau$. In this way, by tuning the noise amplitude —or equivalently the flexibility of the metabolic constraints— the model is able to account for the experimental fact that diversification is observed in variable proportion of the experiments.
@@
Furthermore, as shown in panels (e-g) the model can also explain the empirically-observed variability in both the branching times and the proportion of acetate scavengers after population splitting. Indeed, panel (e) reveals that branching times —estimated as explained in SM Sec. IV— are variable and their distribution depends on the noise amplitude, $\sigma$. One can observe that, in the case with a ”hard” trade-off (i.e. $\sigma=0$), branching occurs almost always at the same time (strongly peaked distribution), but, the standard deviation of the noise $\sigma$ —or equivalently, the amplitude of the trade-off— increases, the distributions of branching times become wider and shift further to the right, corresponding to longer branching times. On the other hand, Fig.2f shows that the distribution of the fraction of acetate scavengers at different times after branching is sharply peaked in the case with a hard trade-off, whereas the case of stochastic trade-off in Fig.2g shows a much broader distribution.
@@
As previously mentioned, the finite widths that emerge in the absence of stochasticity in the trade-off (i.e., for $\sigma=0$) arise from the limited time allowed for the ecological dynamics to stabilize. This is because we assume weak mutations, but not the infinitely-weak mutation limit. As a result, clonal interference and the intrinsic randomness of the evolutionary algorithm may occur (see SM Sec. III). We have computationally verified that the distributions gradually narrow as the time allowed for the ecological dynamics to reach equilibrium is extended (data not shown).
@@ III.2 Branching feasibility and trade-off (in)dependence
Figure 3: Branching feasibility and the fraction of realizations that undergo branching. a) Branching feasibility plot: contour plot that illustrates in phenotypic space the range of local curvatures of the hard trade-off that allow evolutionary branching. This curvature range can also be interpreted (with the stochastic trade-off idea) as the likeliness of having evolutionary branching at each point. b-c) Percentage of realizations (total number, $N_{r}=200$) of the eco-evolutionary model with the stochastic trade-off that undergo evolutionary branching for two different maximum evolutionary times $t_{\text{evo}}^{\text{max}}=2.5\cdot 10^{4}$ and $t_{\text{evo}}^{\text{max}}=5\cdot 10^{4}$ at different points in phenotypic space (that have different curvature range or equivalently different branching feasibility). Points in (b) are those in yellow in panel (a) and $\sigma=10^{-4}$. Points in (c) are those in bourdeaux in panel (a) and $\sigma=3.5\cdot 10^{-4}$. System’s parameters are: $\beta=0.25$ and those specified in SM Table I where the shown $V_{g/a}$ are rescaled with respect to the table value as $V_{a/g}=\gamma V_{a/g}^{\text{table}}$ with $\gamma=4.5\cdot 10^{10}$. Stochastic trade-off parameters are: $\tau=10$ and the others are obtained for each point shown in panel a) as specified in SM Sec. II with $\delta=0.25$. Eco-evolutionary simulations are carried out as described in SM Sec. III.
@@
To this aim, we make use of the formalism proposed by de Mazancourt and Dieckmann [44]. As originally developed, such a framework allows one to analyse geometrically the possible evolutionary outcomes that can occur in the deterministic setting without imposing a specific mathematical function for the trade-off (for the sake of completeness, an overview of the method is described in SM Sec. I). In particular, within this formalism one can construct plots such as that of Fig. 3a, providing information on how independent is evolutionary branching on the specific mathematical form of the hard trade-off. More specifically, it gives at each point in phenotypic space, a measure of the range of possible local curvatures of the hard trade-off function that allows the model to undergo evolutionary branching at such a point, thus quantifying the robustness of branching against changes in the shape of the trade-off function (see SM Sec. I). Regions where the range of possible local curvatures is small (deep blue in Fig. 3a), allow the model to undergo evolutionary branching deterministically only for very specific (fine-tuned) trade-off curves. Instead, in regions where the curvature range is high (light green in Fig. 3a), evolutionary branching occurs deterministically for a larger family of possible trade-off functions (see SM Sec. I for further details).
Now, we would like to see how this (deterministic) formalism could be reinterpreted when we include the idea of stochastic trade-offs. It turns out that using stochastic trade-offs, the previous plots (e.g. Fig.3a) can be interpreted as “evolutionary-branching feasibility plots”, i.e. graphs showing the regions in phenotypic space where it is more likely to see an evolutionary branching event. In other words, regions in phenotypic space where branching is more robust against changes in the specific form of the hard trade-off correspond to regions in phenotypic space where branching is more likely to happen with the stochastic trade-off and vice-versa.
To justify this claim, we employ without loss of generality our eco-evolutionary model with a certain stochastic trade-off, whose average curve produces branching deterministically (i.e. in the $\sigma=0$ limit) at a selected point in phenotypic space (see SM Sec. II for further details on how the stochastic trade-off is built). Then, as illustrated in Fig.3b-c, points where deterministic branching depends weakly on the specific form of the hard trade-off —i.e. points where the range of local curvatures of the hard trade-off that allow branching is larger (brighter colours in Fig.3a)— correspond in the stochastic trade-off case to more realizations that end up experiencing diversification at that point and vice-versa. Note, however, that in this framework we can only work out the relative probabilities of branching. In other words, we can only indicate whether or not one branching condition is relatively more likely compared to another one.
@@
Therefore, interpreting Fig.3a as a branching-feasibility plot, we observe that branching is more likely for low values of $V_{a}$, which agrees with the reported picture of the evolution experiments given in [15]. As shown in Sec.III.3 of the SM, the specific contour pattern depends highly on the conditions of the experiment, that is, the concentration of glucose in the environment supplying the chemostat, $G_{0}$, the dilution rate, $d$, and the proportion of the consumed glucose that gets excreted as acetate, $\beta$. Thus, Fig.3a reveals essentially that the system, under these specific environmental conditions, is rather likely.
@@ III.3 Theoretical predictions for different experimental values
Figure 4: Branching feasibility and the percentage of realizations that undergo branching for different experimental conditions. a-c) Branching-feasibility plots for different values of glucose concentration in the environment supplying the chemostat: $G_{0}=260$ (a), $G_{0}=694$ (b) and $G_{0}=3469$ (c), all of them in $\mu$mol/L. d) Number of realizations that undergo branching at the point depicted in a-c) for the same stochastic trade-off with $(a_{tr},b_{tr},c_{tr},\tau,\sigma)=(88.2,-2.5,0.014,10,6\cdot 10^{-4})$ (see SM for definitions and further information). e-g) Branching-feasibility plots for different values of acetate concentration in the environment supplying the chemostat: $A_{0}=100$ (e), $A_{0}=260$ (f) and $A_{0}=1000$ (g), all of them in $\mu$mol/L. h) Number of realizations that undergo branching at the point depicted in e-g) for the same stochastic trade-off with $(a_{tr},b_{tr},c_{tr},\tau,\sigma)=(164.3,-3.6,0.011,10,5\cdot 10^{-4})$. i-k) Branching-feasibility plots for different values of the dilution rate: $D=0.0015$ (i), $D=0.006$ (j) and $D=0.015$ (k), all of them in 1/min. l) Number of realizations that undergo branching at the point depicted in i-k) for the same stochastic trade-off with $(a_{tr},b_{tr},c_{tr},\tau,\sigma)=(65.5,-5.6,0.088,10,1.2\cdot 10^{-4})$. Total number of realizations for the histograms is $N_{r}=200$. The remaining parameters specified in SM Table I where the shown $V_{g/a}$ are rescaled with respect to the table value as $V_{a/g}=\gamma V_{a/g}^{\text{table}}$ with $\gamma=4.5\cdot 10^{10}$.
@@
In particular, Fig.4a-c clearly illustrates that higher concentrations of glucose in the environment supplying the chemostat ($G_{0}$) generally increase the likelihood of branching. This is further verified in Fig.4d. Similar to Fig.3b-c, we use the same stochastic trade-off parameters for the three different environments whose average curve induces deterministic branching at a selected point. As evident in the histogram, the higher the glucose supply, the greater the number of realizations that undergo evolutionary branching. Moreover, the similarity in the contour pattern shape to that in Fig.3a is attributed to the fact that the proportion of available acetate (i.e. $\beta$ in this case) remains the same.
@@
Finally, it becomes clear from Fig.4i-k that the dilution rate is also a crucial parameter for the emergence of branching. Specifically, higher values of $D$ significantly reduce the likelihood of branching, as illustrated in Fig. 4l.
@@ IV Discussion
Previous work [31] showed, in the context of adaptive dynamics, that a simple eco-evolutionary model can explain in a parsimonious way such a diversification event as the result of evolutionary branching. For this one assumes a specific (hard) trade-off function between the evolving metabolic phenotypes, describing the relative preference for glucose, acetate, or a combination of both. Nevertheless, such an approach is deterministic in nature, i.e., either predicts that branching should be observed or not depending on the chosen trade-off analytical form. Therefore, it is not appropriate to explain the empirically observed stochasticity in the emergence of diversification and the variability both in branching times and in the relative weights of the emerging populations.
Alternatively, in this work, we propose the idea of a stochastic trade-off that stands as an effective representation of the underlying metabolic relations that constrain the evolution of bacteria in a high-dimensional phenotypic space. This high-dimensional space includes factors beyond those directly controlling metabolic preferences. The projection of a given mutational trajectory in such a high-dimensional manifold onto the two-dimensional phenotypic space of glucose-acetate preference does not need to be constrained to a simple curve (as illustrated in Fig. 1). Thus, instead of forcing the evolutionary trajectories in a lower-dimensional metabolic-preference space fixed trade-off curve, stochastic tradeoffs allow the populations to move in a a wider region in such a space, introducing a new source of variability along such a space. In other words, the stochasticity of mutations in a high-dimensional feature space translates into variability that is not strictly constrained to a one-dimensional curve in the space of preferences, but rather to a flexible tradeoff curve. Mathematically, the stochastic trade-off is characterized by a central curve that stands as the average of the different trajectories, the correlation time, $\tau$, that defines how fast the trajectories go back to this average curve and its amplitude, which, for fixed $\tau$, is determined by $\sigma$ and defines the width of the (fuzzy) region where populations can evolve.
Thus, we have built a biologically plausible eco-evolutionary model (see Eqs.1-3), fed with experimentally measured parameters (see SM Table I) and equipped with the mentioned stochastic trade-off. We have shown that such a model is indeed able to account for the variability reported in the experiments (Fig.2). In particular, contrary to existing models it does reproduce: (i) the fact that branching occurs in a large proportion, but not all, of the realizations (Fig.2a-d) (ii) it may happen at considerably different times (Fig.2f) and (iii) the fraction of acetate scavengers at a certain time are variable across experiments (Fig.2g-h).